# TimeCapsule: Solving the Jigsaw Puzzle of Long-Term Time Series Forecasting with Compressed Predictive Representations

### v1

@@ Abstract.
Recent deep learning models for Long-term Time Series Forecasting (LTSF) often emphasize complex, handcrafted designs, while simpler architectures like linear models or MLPs have often outperformed these intricate solutions. In this paper, we revisit and organize the core ideas behind several key techniques, such as redundancy reduction and multi-scale modeling, which are frequently employed in advanced LTSF models. Our goal is to streamline these ideas for more efficient deep learning utilization. To this end, we introduce TimeCapsule, a model built around the principle of high-dimensional information compression that unifies these techniques in a generalized yet simplified framework. Specifically, we model time series as a 3D tensor, incorporating temporal, variate, and level dimensions, and leverage mode production to capture multi-mode dependencies while achieving dimensionality compression. We propose an internal forecast within the compressed representation domain, supported by the Joint-Embedding Predictive Architecture (JEPA), to monitor the learning of predictive representations. Extensive experiments on challenging benchmarks demonstrate the versatility of our method, showing that TimeCapsule can achieve state-of-the-art performance. The code is available for review at: https://anonymous.4open.science/r/TimeCapsule-2BF3.
@@
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Machine learningccs: Information systems Data mining
@@ 1. Introduction
Figure 1. We visualize the ETTh1 dataset as an example to illustrate the need to compress the information contained in different dimensions of MvTS datasets.
With the rapid advancement of deep learning models, Long-term Time Series Forecasting (LTSF) has recently gained prominence. Unlike short-term forecasting, LTSF has traditionally posed significant challenges for classical statistical and machine learning methods, such as VAR, ARIMA, and random forests (Toda and Phillips, 1994; Box and Pierce, 1970; Kane et al., 2014). A series of groundbreaking works (Zhou et al., 2021; Zeng et al., 2023; Nie et al., 2023; Liu et al., 2024b) have been proposed to push the boundaries of this field, addressing limitations and advancing the community. Nevertheless, key questions remain about how best to improve existing models, leaving room for highlighting potential synergies among existing methods and combining their strengths. For instance, Informer suggests that the learned attention map of transformers should be sparse and can be distilled into smaller representations. Models like FiLM (Zhou et al., 2022a) and FEDformer (Zhou et al., 2022b) manipulate time series in compact frequency bases to capture key temporal correlations. These advancements raise several compelling questions: Can these enhancements be conveniently extended to other dimensions? If so, a generalized transformation domain would be required. Is time series data—or the information it contains—inherently compressible? We can intuitively sense this point from two perspectives, as illustrated in Fig. 1.
@@
While these methods individually have provided promising directions, it is shown that different time series may have different underlying preferences (Shao et al., 2024), simplifying and combining these critical principles to build a versatile forecastor is appealing.
Recent discussions in the community have also drawn attention to the surprisingly strong performance of linear models when compared to transformer-based architectures in LTSF tasks (Zeng et al., 2023). While transformers (Vaswani, 2017) have set milestones across domains like Natural Language Processing (NLP) (Wang et al., 2022) and Computer Vision (CV) (Yuan et al., 2021), their effectiveness in LTSF remains inconsistent, often falling short of simpler linear models (Zeng et al., 2023). In this work, we propose that an effective architecture for LTSF should simply consist of two stages: predictive representation learning and generalized linear projection, in which the latter serves as an accurate predictor to learn generalized linear dependencies for the forecasting, while the former extracts abstract and informative representation from distinct datasets to improve forecasters’ generality. Therefore, we propose TimeCapsule, a novel model employing a Chaining Bits Back with Asymmetric Numeral Systems (BB-ANS) (Townsend et al., 2019) like-architecture, which adopts weak transformer-based blocks as the encoder, while MLP-based blocks as the decoder, striking a balance between powerful representation learning and computational simplicity. Technically, our model is driven by three key principles:
Multi-level Modeling: Multi-scale modeling stands out as an effective paradigm for improving LSTF performance (Ferreira et al., 2006). Existing methods incorporate multiresolution analysis by up/downsampling through moving average and convolutional pooling layers (Challu et al., 2023; Wang et al., 2024), or designing hierarchical structures (Liu et al., 2021; Zhang and Yan, 2023) to aggregate multi-scale features. Besides, time series decomposition is also a traditional and commonly used strategy to improve LTSF performance (Oreshkin et al., 2019; Zhou et al., 2022b; Wu et al., 2021). In avoidance of complexity, we propose adding an extra dimension called level to the original time series, allowing the model to learn multi-level features within the representation space. This approach generalizes multi-scale learning and time series decomposition, in the meantime, making the learning process model-independent.
@@
Our contributions can be summarized as follows:
@@ 2.1. Long-term Time Series modeling
The first group (e.g., Autoformer (Wu et al., 2021), N-BEATs (Oreshkin et al., 2019), Pyraformer (Liu et al., 2021), N-Hits (Challu et al., 2023), Crossformer (Zhang and Yan, 2023), TimeMixer (Wang et al., 2024)) focuses on multi-level modeling, which incorporates multiresolution/multi-scale analysis and series decomposition within the model. These techniques enable the model to learn both coarse and fine-grained features within time series, facilitating the capture of hierarchical temporal patterns. The second group (e.g., Informer (Zhou et al., 2021), FEDformer (Zhou et al., 2022b), PatchTST (Nie et al., 2023)), leverages information redundancy to filter and extract high-energy temporal correlations, or models time series in the form of temporal patches, avoiding the inclusion of finer, noisier information in the forecasting process, improving the model’s robustness and prediction accuracy., These advancements lead to progressive improvements in both effectiveness and efficiency. The third group (e.g., Crossformer, iTransformer (Liu et al., 2024b)) explores the impact of attention applied across various dimensions of time series, offering novel perspectives for time series correlation extraction. This group demonstrates that capturing dependencies across different modes can significantly enhance forecasting performance, by providing a richer understanding of temporal and variable relationships. The fourth group (e.g., DLinear (Zeng et al., 2023), N-BEATs, N-Hits, TsMixer (Chen et al., 2023), TimeMixer) emphasizes the importance of generalized linear dependency modeling and capitalizes on linear or MLP-based architectures to establish highly effective forecasters. This validates the crucial role of learning the appropriate coefficients to combine the captured base components of time series and the dependencies they contain.
@@ 3. Preliminaries
where $\textbf{M}\in\mathbb{R}^{m\times n_{3}}$ is the transform factor. This production can be generalized to any mode of any tensor, leading to the definition of mode-$k$ product.
@@ 4.2.1. Mode Specific Multi-head Self-Attention (MoMSA)
where ×$\times_{2}$ denotes the mode-2 product, which represents a matrix multiplication along the second dimension of a 3D tensor. T-MSA denotes the vanilla multi-head self-attention (Vaswani, 2017) applied to the T (temporal) dimention. It is notable that with the setting of $m\leq v_{a}$, the dimension can be compressed to an arbitrary length, resulting in a compressed attention map.
When applying this procedure to the real case, we shall obtain $\mathcal{X}_{1}\in\mathbb{R}^{v\times t_{c}\times 1}$ from the first pipe as shown in the Fig.3. However, it is risky to make such a compression on the original MvTS information. To address this, we take a number of protective measures within our TransBlock. In particular, before applying MoMSA, we project the information into the embedding space $\mathbb{R}^{t_{x}\times d}$ and introduce Gaussian noise at the start of the block, which may improve robustness during compression (channel coding). Furthermore, we set the transform factor as the product of two matrices
@@ 4.2.3. Representation Prediction with JEPA Loss
In addition to the time-domain prediction, the structure of TimeCapsule introduces inner predictions for the compressed representation. We hypothesize a potential gap (e.g., inter-space distribution shift (Dish-TS)) between the historical and future series. We argue that the use of reversible instance normalization (RevIn) (Kim et al., 2021) does not affect the stationarity of the time series. The primary objective of RevIn is to mitigate distribution shifts between the training and testing data. However, because the statistics it uses are calculated along the temporal dimension, the shape and stochastic properties of the series remain invariant.
A key challenge is how to track and validate this inner prediction. To address this, we employ the Joint-Embedding Predictive Architecture (JEPA), which not only offers a measure to evaluate the distance between observations and predictions in non-stationary, time-varying processes but also enables efficient contrastive loss in the representation space. This helps the internal prediction converge, and in some cases, may improve the real prediction as well.
To implement JEPA, another encoder is required to convert the target into the same representation space as the input. In line with the strategies used in I-JEPA and V-JEPA, we continue to obtain the target encoder by applying the Exponential Moving Average (EMA) of the input encoder. However, due to the nature of LTSF, how to deal with the inconsistency of the sequence length between the input and output is problematic. We disentangle this problem by preprocessing the target sequence to match the input length as follows,
i.e., in cases where $t_{y}<t_{x}$, we pad Y with zeros; and when $t_{y}$ is larger, we can reasonably consider the whole series as a carrier of information, and that the correlations between nearby timestamps are stronger. Under this assumption, we split the sequence Y into sub-sequences with length $t_{x}$, then apply EMA to these chunks to form integrated time information. This enables an efficient computation of the predictive representation loss:
$\text{Loss}(\text{Enc}_{x}(\text{X}),~{}sg(\text{Enc}_{y}(\text{Y}))$
5. Experiment
5.1. Forecasting Results
5.1.1. Datasets and Baselines.
We utilize datasets from various domains, including electricity (ETTh1, ETTh2, ETTm1, ETTm2, Electricity), environment (Weather), energy (Solar-Energy), transportation (PEMS04, Traffic), and health (ILI). The forecasting methods we compare against include iTransformer (Liu et al., 2024b), TimeMixer (Wang et al., 2024), PatchTST (Nie et al., 2023), Crossformer (Zhang and Yan, 2023), DLinear (Zeng et al., 2023), TimesNet (Wu et al., 2022), FEDformer (Zhou et al., 2022b), and Informer (Zhou et al., 2021).
5.1.2. Main Configurations.
5.1.3. Main Results.
Table 1. Full results of multivariate forecasting. For TimeCapsule, we use upperscript $\dagger$ to denote the employment of JEPA training. The lookback length $T=96$ and prediction lengths $S\in\{24,36,48,60\}$ for ILI, $S\in\{96,192,336,720\}$ and fixed lookback length $T=512$ for others; For other models, lookback lengths are searched for the best performance.
@@ 5.2. Model Analysis
Table 2. Ablation on the information compensation design. We remove the residual information connection or replace $\mathcal{B}$ by the original information $\mathcal{X}$. The resulting performance variations are then presented in terms of mean squared error (MSE).
@@
We compare TimeCapsule with other deep learning techniques designed to handle non-stationarity in time series. The results show that TimeCapsule outperforms other models in trend prediction. Interestingly, for the ETTh1 dataset, MLP achieves similar performance to TimeCapsule, though its trend predictions are not as sound as TimeCapsule. This difference can be attributed to two factors. First, While not always accurate, TimeCapsule’s implicit internal forecasting mechanism allows it to inherently predict the non-stationary trend, alongside other models designed for such tasks. Second, the use of RevIn helps TimeCapsule predict at the correct scales. More visualizations are provided in Appendix G.
@@
In contrast, the results for the PEMS04 and Electricity datasets show a growing trend in JEPA loss. To further explore this, we compare forecasting accuracy with their original values at the bottom part of Fig. 6. From this, we can empirically conclude that in many cases, regularizing the optimization with the inductive bias introduced by JEPA can be beneficial. Incorporating JEPA loss into the training process can encourage the forecaster to converge to a favorable minimum before stepping into other directions. However, the effect of attaching such a reinforcement becomes negligible when the descent directions of two losses align or when a flat energy plane is reached. This finding may provide greater flexibility in the use of JEPA for optimization guidance in predictive representation learning.
@@
We investigate the asymmetric structure of TimeCapsule by varying the compression dimensions to figure out which part—transformer-based encoder or MLP-based decoder—plays a more crucial role. By default, the compression dimensions are set to {4, 8, 4}. Then we test two alternative configurations: (a) {4, 1, 4}, which removes the level embedding effect, and (b) {1, 1, 1}, which reduces the entire model to a pure MLP. The results in Table 3 reveal that for datasets with higher sampling frequencies, such as PEMS04, multi-level modeling and capturing multi-mode dependencies show significant benefits. In contrast, for datasets like weather and electricity, a simpler MLP structure appears to be sufficient for long-term forecasting (see Appendix C.5 for results of more datasets). These observations, along with the comprehensive results in Table 1, suggest that designing a universal, efficient model for diverse datasets is challenging, often leading to inefficient module allocation. Furthermore, a forecaster with strong generalized linear modeling capacity can manage most cases effectively and is more likely to be the key for LTSF.

### v3

@@ Abstract.
Recent deep learning models for Long-term Time Series Forecasting (LTSF) often emphasize complex, handcrafted designs, while simpler architectures like linear models or MLPs have often outperformed these intricate solutions. In this paper, we revisit and organize the core ideas behind several key techniques, such as redundancy reduction and multi-scale modeling, which are frequently employed in advanced LTSF models. Our goal is to streamline these ideas for more efficient deep learning utilization. To this end, we introduce TimeCapsule, a model built around the principle of high-dimensional information compression that unifies these techniques in a generalized yet simplified framework. Specifically, we model time series as a 3D tensor, incorporating temporal, variate, and level dimensions, and leverage mode production to capture multi-mode dependencies while achieving dimensionality compression. We propose an internal forecast within the compressed representation domain, supported by the Joint-Embedding Predictive Architecture (JEPA), to monitor the learning of predictive representations. Extensive experiments on challenging benchmarks demonstrate the versatility of our method, showing that TimeCapsule can achieve state-of-the-art performance. The code is available at: https://github.com/Luoauoa/TimeCapsule.git.
@@
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3711896.3737157isbn: 979-8-4007-1454-2/2025/08ccs: Computing methodologies Machine learningccs: Information systems Data mining
@@ 1. Introduction
With the rapid advancement of deep learning models, Long-term Time Series Forecasting (LTSF) has recently gained prominence. Unlike short-term forecasting, LTSF has traditionally posed significant challenges for classical statistical and machine learning methods, such as VAR, ARIMA, and random forests (Toda and Phillips, 1994; Box and Pierce, 1970; Kane et al., 2014). A series of groundbreaking works (Zhou et al., 2021; Zeng et al., 2023; Nie et al., 2023; Liu et al., 2024b) have been proposed to push the boundaries of this field, addressing limitations and advancing the community. Nevertheless, questions remain about how best to improve existing models, leaving room for highlighting potential synergies among existing methods and combining their strengths. For instance, Informer(Zhou et al., 2021) suggests that the learned attention map of transformers should be sparse and can be distilled into smaller representations. Models like FiLM (Zhou et al., 2022a) and FEDformer (Zhou et al., 2022b) manipulate time series in compact frequency bases to capture key temporal correlations. These advancements raise several compelling questions: Can these enhancements be conveniently extended to other dimensions? If so, a generalized transformation domain would be required. Moreover, is time series data—or the information it contains—inherently compressible? We can intuitively sense this point from two perspectives, as illustrated in Fig. 1.
@@
Figure 1. We visualize the ETTh1 dataset as an example to illustrate the need to compress the information in different MvTS datasets’ dimensions.
@@
While these methods individually have provided promising directions, it is shown that different time series may have different underlying preferences (Shao et al., 2024). Simplifying and combining these critical principles to build a versatile forecastor that can identify the intrinsic property of each dataset is appealing.
Recent discussions in the community have also drawn attention to the surprisingly strong performance of linear models when compared to transformer-based architectures in LTSF tasks. It turns out that the effectiveness of transformers (Wang et al., 2022; Yuan et al., 2021) in LTSF remains inconsistent, often falling short of simpler linear models (Zeng et al., 2023). In this work, we propose that an effective architecture for LTSF should consist of two stages: predictive representation learning and generalized linear projection, in which the latter serves as an accurate predictor to learn generalized linear dependencies for the forecasting, while the former extracts abstract and informative representation from distinct datasets to improve forecasters’ generality. Therefore, we propose TimeCapsule, a novel model employing a Chaining Bits Back with Asymmetric Numeral Systems (BB-ANS) (Townsend et al., 2019) like-architecture, which adopts weak transformer-based blocks as the encoder, while MLP-based blocks as the decoder, striking a balance between powerful representation learning and computational simplicity. Technically, our model is driven by three principles:
Multi-level Modeling: Multi-scale modeling stands out as an effective paradigm for improving LSTF performance (Ferreira et al., 2006). Existing methods incorporate multiresolution analysis by up/downsampling through moving average and convolutional pooling layers (Challu et al., 2023; Wang et al., 2024), or designing hierarchical structures (Liu et al., 2021; Zhang and Yan, 2023) to aggregate multi-scale features. Besides, time series decomposition is also a traditional and commonly used strategy to improve LTSF performance (Oreshkin et al., 2019; Zhou et al., 2022b; Wu et al., 2021). In avoidance of complexity, we propose adding an extra dimension called level to the original time series, allowing the model to learn multi-level features within the representation space. This approach generalizes multi-scale learning and time series decomposition, making the learning process model-independent.
@@
Through this framework, we hope the neural network not only purifies redundant information but also learns to decompose, analyze, and recreate the underlying mechanisms of diverse time series. Our contributions can be summarized as follows:
@@ 2.1. Long-term Time Series modeling
The first group (e.g., Autoformer (Wu et al., 2021), N-BEATs (Oreshkin et al., 2019), Pyraformer (Liu et al., 2021), N-Hits (Challu et al., 2023), Crossformer (Zhang and Yan, 2023), TimeMixer (Wang et al., 2024)) focuses on multi-level modeling, which incorporates multiresolution/multi-scale analysis and series decomposition within the model. These techniques enable the model to learn both coarse and fine-grained features within time series, facilitating the capture of hierarchical temporal patterns. The second group (e.g., Informer (Zhou et al., 2021), FEDformer (Zhou et al., 2022b), PatchTST (Nie et al., 2023)), leverages information redundancy to filter and extract high-energy temporal correlations, or models time series in the form of temporal patches, avoiding the inclusion of finer, noisier information in the forecasting process, improving the model’s robustness and prediction accuracy., These advancements lead to progressive improvements in both effectiveness and efficiency. The third group (e.g., Crossformer, iTransformer (Liu et al., 2024b)) explores the impact of attention applied across various dimensions of time series, offering novel perspectives for time series correlation extraction. This group demonstrates that capturing dependencies across different modes can significantly enhance forecasting performance by providing a richer understanding of temporal and variable relationships. The fourth group (e.g., DLinear (Zeng et al., 2023), N-BEATs, N-Hits, TsMixer (Chen et al., 2023), TimeMixer) emphasizes the importance of generalized linear dependency modeling and capitalizes on linear or MLP-based architectures to establish highly effective forecasters. This validates the crucial role of learning the appropriate coefficients to combine the captured base components of time series and the dependencies they contain.
@@ 3. Preliminaries
where $\textbf{M}\in\mathbb{R}^{m\times n_{3}}$ is the transform factor. This production can be generalized to any mode of any tensor, leading to the definition of the mode-$k$ product.
@@ 4.2.1. Mode Specific Multi-head Self-Attention (MoMSA)
where ×$\times_{2}$ denotes the mode-2 product, which represents a matrix multiplication along the second dimension of a 3D tensor. T-MSA denotes the vanilla multi-head self-attention (vaswani2017attention) applied to the T (temporal) dimention. It is notable that with the setting of $m\leq v_{a}$, the dimension can be compressed to an arbitrary length, resulting in a compressed attention map.
When applying this procedure to the real case, we shall obtain $\mathcal{X}_{1}\in\mathbb{R}^{v\times t_{c}\times 1}$ from the first pipe as shown in Fig.  ref fig:2. However, it is risky to make such a compression on the original MvTS information. To address this, we take a number of protective measures within our TransBlock. In particular, before applying MoMSA, we project the information into the embedding space $\mathbb{R}^{t_{x}\times d}$ and introduce Gaussian noise at the start of the block, which may improve robustness during compression (channel coding). Furthermore, we set the transform factor as the product of two matrices
@@ 4.2.3. Representation Prediction with JEPA Loss
In addition to the time-domain prediction, the structure of TimeCapsule introduces inner predictions for the compressed representation. We hypothesize a potential gap (e.g., inter-space distribution shift (Dish-TS)) between the historical and future series. We argue that the use of reversible instance normalization (RevIn) (Kim et al., 2021) does not affect the stationarity of the time series. The primary objective of RevIn is to mitigate the distribution shifts between the training and testing data. However, because the statistics it uses are calculated along the temporal dimension, the shape and stochastic properties of the series remain invariant.
A key challenge is how to track and validate this inner prediction. To address this, we employ the Joint-Embedding Predictive Architecture (JEPA), which not only offers a measure to evaluate the distance between observations and predictions in non-stationary, time-varying processes but also enables efficient contrastive loss in the representation space. This helps the internal prediction to converge and, in some cases, may improve the real prediction as well.
Table 1. Full results of multivariate forecasting. For TimeCapsule, we use superscript $\dagger$ to denote the employment of JEPA training. The lookback length $T=96$ and prediction lengths $S\in\{24,36,48,60\}$ for ILI, $S\in\{96,192,336,720\}$ and fixed lookback length $T=512$ for others; For other models, lookback lengths are searched for the best performance.
@@
To implement JEPA, another encoder is required to convert the target into the same representation space as the input. In line with the strategies used in I-JEPA and V-JEPA, we continue to obtain the target encoder by applying the Exponential Moving Average (EMA) of the input encoder. However, due to the nature of LTSF, it is problematic to deal with the inconsistency of the sequence length between the input and output. We disentangle this problem by preprocessing the target sequence to match the input length as follows,
i.e., in cases where $t_{y}<t_{x}$, we pad Y with zeros; and when $t_{y}$ is larger, we can reasonably consider the whole series as a carrier of information, and the correlations between nearby timestamps are stronger. Under this assumption, we split the sequence Y into subsequences of length $t_{x}$, then apply EMA to these chunks to form integrated time information. This enables an efficient computation of the predictive representation loss:
$\text{Loss}(\text{Enc}_{x}(\text{X}),~{}sg(\text{Enc}_{y}(\text{Y}))$
5. Experiment
5.1. Forecasting Results
5.1.1. Datasets and Baselines.
We utilize datasets from various domains, including electricity (ETTh1, ETTh2, ETTm1, ETTm2, Electricity), environment (Weather), energy (Solar-Energy), transportation (PEMS04, Traffic), and health (ILI). The forecasting methods against which we compare include iTransformer (Liu et al., 2024b), TimeMixer (Wang et al., 2024), PatchTST (Nie et al., 2023), Crossformer (Zhang and Yan, 2023), DLinear (Zeng et al., 2023), TimesNet (Wu et al., 2022), FEDformer (Zhou et al., 2022b) and Informer (Zhou et al., 2021).
5.1.2. Main Configurations.
5.1.3. Main Results.
@@ 5.2. Model Analysis
Table 2. Ablation on the information compensation design. We remove the residual information connection or replace $\mathcal{B}$ with the original information $\mathcal{X}$. The resulting performance variations are then presented in terms of mean squared error (MSE).
@@
Figure 4. We visualize ETTh1 dataset as a typical example to illustrate 512-96 predictions of different components (trend, stationary, deviation) of time series. All results are obtained under the recommended settings of the source code.
@@
We compare TimeCapsule with other deep learning techniques designed to handle non-stationarity in time series. The results show that TimeCapsule outperforms other models in trend prediction. Interestingly, for the ETTh1 dataset, MLP achieves similar performance to TimeCapsule, though its trend predictions are not as sound as TimeCapsule. This difference can be attributed to two factors. First, While not always accurate, TimeCapsule’s implicit internal forecasting mechanism allows it to inherently predict the non-stationary trend, alongside other models designed for such tasks. Second, the use of RevIn helps TimeCapsule predict at the correct scales.
In contrast, the results for the PEMS04 and Electricity datasets show a growing trend in JEPA loss. To further explore this, we compare forecasting accuracy with their original values at the bottom part of Fig. 5. From this, we can empirically conclude that in many cases, regularizing the optimization with the inductive bias introduced by JEPA can be beneficial. Incorporating JEPA loss into the training process can encourage the forecaster to converge on a favorable minimum before stepping into other directions. However, the effect of attaching such a reinforcement becomes negligible when the descent directions of two losses align or when a flat energy plane is reached. This finding may provide greater flexibility in the use of JEPA for optimization guidance in predictive representation learning.
@@
We investigate the asymmetric structure of TimeCapsule by varying the compression dimensions to figure out which part—transformer-based encoder or MLP-based decoder—plays a more crucial role. By default, the compression dimensions are set to {4, 8, 4}. Then we test two alternative configurations: (a) {4, 1, 4}, which removes the level embedding effect, and (b) {1, 1, 1}, which reduces the entire model to a pure MLP. The results in Table 3 reveal that for datasets with higher sampling frequencies, such as PEMS04, multi-level modeling and capturing multi-mode dependencies show significant benefits. In contrast, for datasets like weather and electricity, a simpler MLP structure appears to be sufficient for long-term forecasting. These observations, along with the comprehensive results in Table 1, suggest that designing a universal, efficient model for diverse datasets is challenging, often leading to inefficient module allocation. Furthermore, a forecaster with strong generalized linear modeling capacity can manage most cases effectively and is more likely to be the key to LTSF.